The accompanying data were obtained from the financial aid office of a certain university Receiving Financial Aid 4,226 1,879 6,105 Not Receiving Financial Aid 3,892 731 4,623 Total 8,118 2,610 10,728 Undergraduates Graduates Total Let A be the event that a student selected at random from this university is an undergraduate student, and let B be the event that a student selected at random is receiving financial aid (a) Find each of the following probabilities. (Round your answers to three decimal places.) P(A)757 P(B) = .568 P(A n B)394 P(B | A) .521 P(B I Ac)717
Answer #1

a) P(B) = 6105/10728 = 0.569

b) P(B | Ac) = P(Receiving financial aid | Graduates) = 1879/2610 = 0.720